In the first half of a multi-year temperature record at some location, the mean temperature and uncertainty are 5.5 and 0.06 °C, respectively. For the second half of the record, the mean and uncertainty are 5.7 and 0.06 °C, respectively Explain in the box below using numerical values and associated units whether there has been a statistically significant change in annual mean temperature over these two periods HINT: Determine whether the difference between the two 25-year period means is more than twice the uncertainty in the difference (i.e. 95% confidence).
Answer #1

here as std deviation for both the temperature records are same ; hence pooled std deviation for difference in temperature will remain same,

here estimated mean difference =5.5-5.7=-0.2

therefore 95% confidence interval for mean difference =estimated mean difference -/+2*std deviation

=-0.2 -/+ 2*0.06 =-0.32 to -0.08

as confidence interval values are all below 0 which represent that there is statistically significant difference change in annual mean temperature over these two periods